OpenStudy (anonymous):

I need someone SMART in algebra 2 2In(x)+3In(2)=5 i have to evaluate and check it

Nnesha (nnesha):

quotient rule\[\large\rm ln y - ln x = \ln \frac{ x }{ y}\] subtraction to division product rule \[\large \rm ln x + \ln y = \ln ( x \times y )\] addition ----> multiplication power rule \[\large\rm ln x^y = y \ln x\] first apply power property

OpenStudy (anonymous):

i got this far which is not much: Inx^2+3In(2)=5

Nnesha (nnesha):

\[\huge\rm Inx^2+\color{Red}{3}In(2)=5\] one more step

OpenStudy (anonymous):

ok

Nnesha (nnesha):

any number or variable at front of log or ln wll become exponent